# Bringing memory to Boolean networks: a unifying framework

### v1

@@
Abstract. Boolean networks are extensively applied as models of complex dynamical systems, aiming at capturing essential features related to causality and synchronicity of the state changes of components along time. Dynamics of Boolean networks result from the application of their Boolean map according to a so-called update mode, specifying the possible transitions between network configurations. In this paper, we explore update modes that possess a memory on past configurations, and provide a generic framework to define them. We show that recently introduced modes such as the most permissive and interval modes can be naturally expressed in this framework. We propose novel update modes, the history-based and trapping modes, and provide a comprehensive comparison between them. Furthermore, we show that trapping dynamics, which further generalize the most permissive mode, correspond to a rich class of networks related to transitive dynamics and encompassing commutative networks. Finally, we provide a thorough characterization of the structure of minimal and principal trapspaces, bringing a combinatorial and algebraic understanding of these objects.
@@ 1 Introduction
A large amount of theoretical work underlined the importance of the scheduling of the update of states of components for the dynamics that can be generated, the so-called update mode. One natural mode is the parallel (or synchronous) mode, where all components are updated simultaneously and instantaneously, generating deterministic dynamics. Another classical update mode is the (general) asynchronous mode: any subset of components can be updated simultaneously, potentially leading to non-deterministic dynamics. Any trajectory, i.e., succession of configurations, computed with the parallel mode can also be computed using the asynchronous mode. The converse is, in general, false. Update modes also impact the so-called limit configurations, that are the configurations from which any reachable configuration can return to them. Limit configurations constitute the attractors of dynamics and are a prime subject of study in the literature, including their robustness to update modes [2, 1].
When employing BNs as models of complex dynamical systems, such as gene regulation within cells, update modes aim at reflecting hypotheses related to the duration, speed, probability, and other quantitative features of the transitions. For instance, asynchronous modes can be motivated by the fact that genes can require different amounts of proteins to become active.
@@
We propose a unifying framework to define such update modes by characterizing the trajectories they generate. This allows to account for a sequence of configurations that have been computed since an initial one, and compute elongation of the trajectory based on this sequence. We notably introduce in Sect. 2 two elementary update modes using memory: the history-based update mode, where the next configuration is computed from the current one and one component has been updated according to any configuration in the past; and the trapping update mode, where the next configuration is computed from any configuration of the past, and one component has been updated according to any configuration in the past. Intuitively, the history-based mode extends the asynchronous by allowing to take any configuration of the past, instead of the latter only, to compute the next state of the component. Then, the trapping mode extends the history-based by allowing to take any configuration of the past as the base next configuration. This somehow allows for jumping back in time, while having the knowledge of the sequence of configurations before the jump.
@@
One can naturally derive that the asynchronous trajectories are a subset of history-based trajectories, being themselves a subset of trapping trajectories. In this paper, we establish simulation and weak simulation relationships with the interval, cuttable, and most permissive update modes (Sect. 3). In particular, we demonstrate that cuttable and history-based modes are incomparable, while both weakly simulating the interval, and both being weakly simulated by the most permissive, which in turns, is weakly simulated by the trapping mode. We provide a similar hierarchy by considering inclusion of trajectories (simulation). We also show that the trapping mode is weakly bisimilar to a subcube-based update mode. The (smallest) subcube enclosing the configurations of the trajectory is the set of configurations that can be composed of any state of the components in the past. In the subcube-based mode, the trajectory can be extended with any of such configurations, where the state of one component has been updated according to any configuration of this subcube.
@@
Another important result of Sect. 3 is that most permissive and trapping modes coincide on the reachability of configurations in minimal trapspaces. Thus, both dynamics have the same limit configurations, and the same limit configurations reachable from any configuration.
Finally, in Sect. 4, we further explore combinatorics of the trapping mode and show that trapping dynamics form a rich class of graphs. In particular, they correspond to general asynchronous dynamics being transitive, and encompass commutative networks. The principal trapspace of a configuration (for a given BN) is the smallest trapspace that contains that configuration. All trajectories in the update modes considered in this paper are bounded by the principal trapspace of their starting point; in fact, the trapping update mode can visit the entire principal trapspace. We then investigate the structure of principal trapspaces. Clearly, every subcube can be a principal trapspace, however the collection of all principal trapspaces of a BN is highly structured (we call it pre-principal in the sequel; see Lemma 4.13). Moreover, the collection of all trapspaces of a BN is also highly structured (roughly speaking, it is closed under union and intersection; see Lemma 4.14). In Theorem 4.9 we show that two BNs have the same trapspaces if and only if they have the same principal trapspaces. In other words, the collection of principal trapspaces of a BN completely characterises the collection of all its trapspaces. We then strengthen this result and show in Theorem 4.11 that there is a three-way equivalence between: BNs with transitive general asynchronous graphs (which we call trapping networks), collections of principal trapspaces of BNs, and collections of trapspaces of BNs. We finally consider minimal trapspaces, which form a subset of principal trapspaces. Unlike principal trapspaces, the collection of minimal trapspaces does not determine the collection of all trapspaces. However, we are able to characterise collections of minimal trapspaces of BNs and to obtain results analogous to Theorems 4.9 and 4.11.
@@
The paper is structured as follows. Sect. 2 introduces a generic framework to express memory-based update modes which is employed to define novel trapping, history-based, and subcube-based update modes, and give equivalent definitions of asynchronous, most permissive, interval and cuttable update modes. Then, Sect. 3 provides a simulation and weak-simulation hierarchy between those updates. Focusing on trapping and most permissive update modes, Sect. 4 brings combinatorial characterization of trapping dynamics and of collections of trapspaces and minimal trapspaces. Finally, Sect. 5 discusses the contributions of the paper.
@@
We denote the Boolean set by $\mathbb{B}=\{0,1\}$ and for any positive integer $n$, we denote $[n]=\{1,\dots,n\}$. A configuration is $x=(x_{1},\dots,x_{n})\in\mathbb{B}^{n}$. For any $S\subseteq[n]$, we denote $x_{S}=(x_{s}:s\in S)$ and we use the notation $x=(x_{S},x_{-S})$. We shall identify an element $i\in[n]$ with the corresponding singleton $\{i\}$, so that $x=(x_{i},x_{-i})$ for instance. For any two configurations $x,y\in\mathbb{B}^{n}$, we denote the set of positions where they differ by $\Delta(x,y)=\{i\in[n]:x_{i}\neq y_{i}\}$ and their Hamming distance by $d_{\mathrm{H}}(x,y)=|\Delta(x,y)|$.
@@
A subcube of $\mathbb{B}^{n}$ is any $X\subseteq\mathbb{B}^{n}$ such that there exist two disjoint sets of positions $S,T\subseteq[n]$ with $X=\{x\in\mathbb{B}^{n}:x_{S}=0,x_{T}=1\}$. For any set $A\subseteq\mathbb{B}^{n}$, the principal subcube of $A$, denoted by $[A]$, is the smallest subcube containing $A$. If $A=\{a_{1},\dots,a_{k}\}$, we also denote $[A]=[a_{1},\dots,a_{k}]$. If $X$ is a subcube and $x\in X$, then there is a unique $y\in X$ such that $X=[x,y]$; we refer to $y$ as the opposite of $x$ in $X$, and we denote it by $y=X-x$. We denote the set of subcubes of $\mathbb{B}^{n}$ as $\mathrm{S}(n)$ and the set of all collections of subcubes of $\mathbb{B}^{n}$ as $\mathrm{A}(n)=2^{\mathrm{S}(n)}$.
@@
A Boolean network (BN) of dimension $n$ is a mapping $f:\mathbb{B}^{n}\to\mathbb{B}^{n}$. We denote the set of BNs of dimension $n$ as $\mathrm{F}(n)$. Any BN $f\in\mathrm{F}(n)$ can be viewed as $f=(f_{1},\dots,f_{n})$ where $f_{i}(x)=f(x)_{i}$ for all $i\in[n]$. For any $S\subseteq[n]$ and any $f\in\mathrm{F}(n)$, the update of $S$ according to $f$ is represented by the BN $f^{(S)}\in\mathrm{F}(n)$ where
$f^{(S)}(x)=(f_{S}(x),x_{-S}).$
In particular, $f^{([n])}=f$ and $f^{(\emptyset)}=\mathrm{id}$. We note the distinction between the update $f^{(i)}$ (given by $f^{(i)}(x)=(f_{i}(x),x_{-i})$) and the power $f^{i}=f\circ f\circ\dots\circ f$ ($i$ terms). We can then compose those updates, so that if $S_{1},\dots,S_{k}\subseteq[n]$, we obtain
$f^{(S_{1},\dots,S_{k})}=f^{(S_{k})}\circ f^{(S_{k-1})}\circ\dots\circ f^{(S_{1})}.$
@@
A trapspace of $f\in\mathrm{F}(n)$ is a subcube $X\subseteq\mathbb{B}^{n}$ such that $f(X)\subseteq X$. The collection of all trapspaces of $f$, denoted by $\mathcal{T}(f)$, is closed under intersection. Then for any $x\in\mathbb{B}^{n}$,there is a smallest trapspace of $f$ that contains $x$, which we shall refer to as the principal trapspace of $x$ (with respect to $f$). For the sake of simplicity, we denote it by $T_{f}(x)$. The principal trapspace $T_{f}(x)$ can be recursively computed as follows: let $T_{0}=\{x\}$ and $T_{i}=[T_{i-1}\cup f(T_{i-1})]$, then $T_{n}=T_{f}(x)$. In particular, $[x,f(x)]\subseteq T_{f}(x)$. The collection of all principal trapspaces of $f$ is denoted by $\mathcal{P}(f)$. A trapspace $T$ is minimal if there is no trapspace $T^{\prime}$ with $T^{\prime}\subset T$. Clearly, any minimal trapspace is principal, but the converse does not necessarily hold. The collection of minimal trapspaces of $f$ is denoted by $\mathcal{M}(f)$. Minimal trapspaces of BNs have been notably studied for their relation with limit (ultimately periodic) configurations: each minimal trapspace necessarily contains at least one limit configuration [11, 14].
A (directed) graph is $\Gamma=(V,E)$, where $V$ is the set of vertices and $E\subseteq V^{2}$ is the set of edges. A graph $\Gamma$ is reflexive if for all $v\in V$, $(v,v)\in E$; $\Gamma$ is symmetric if $(u,v)\in E$ implies $(v,u)\in E$ for all $u,v\in V$; and $\Gamma$ is transitive if $(u,v),(v,w)\in E$ implies $(u,w)\in E$ for all $u,v,w\in E$. The out-neighbourhood of a vertex $v$ is $N^{out}(\Gamma;v)=\{u\in V:(v,u)\in E\}$.
@@
The asynchronous graph of a BN $f\in\mathrm{F}(n)$ is the graph $\mathtt{A}(f)=(V,E)$ where $V=\mathbb{B}^{n}$ and
$E=\{(x,f^{(i)}(x)):x\in\mathbb{B}^{n},i\in[n]\}.$
@@
Note that in most literature, one removes the loops $(x,x)$ from the asynchronous graph, that occur every time $f_{i}(x)=x_{i}$, yet we shall keep those loops instead in our definition. However, when drawing the asynchronous graph, we shall not display the loops and instead draw the underlying hypercube with thin black lines and the arcs of the graph with thick blue arrows. See below for an example of a BN, for which we give the asynchronous graph and all the trapspaces.
@@
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
We then have $\mathcal{T}(f)=\{A,B,C,D,E,F,G,H\}$, $\mathcal{P}(f)=\{A,B,C,D,E,F\}$ and $\mathcal{M}(f)=\{D,E,F\}$ with
$T_{f}(000)=T_{f}(010)$ $=$ $A$ =$=\{x:x_{3}=0\}$
$T_{f}(111)$ $=$ $B$ =$=\{x:x_{12}=11\}$
$T_{f}(011)=T_{f}(001)$ $=$ $C$ =$=\{x\}$
$T_{f}(110)$ $=$ $D$ =$=\{x:x_{12}=11,x_{3}=0\}$
$T_{f}(100)$ $=$ $E$ =$=\{x:x_{1}=1,x_{23}=00\}$
$T_{f}(101)$ $=$ $F$ =$=\{x:x_{13}=11,x_{2}=0\}$
$G$ =$=\{x:x_{1}=1,x_{3}=0\}$
$H$ =$=\{x:x_{1}=1\}.$
@@ 2 Update modes for dynamics with memory
A (fully asynchronous) trajectory is a sequence $(x^{0},\dots,x^{l})\in(\mathbb{B}^{n})^{*}$ such that for all $1\leq a\leq l$, there exists $t\in[x^{0},\dots,x^{a-1}]$ such that $d_{\mathrm{H}}(t,x^{a})\leq 1$. A prefix of the trajectory $(x^{0},\dots,x^{l})$ is any $(x^{0},\dots,x^{a})$ for $0\leq a\leq l$, or the empty sequence. An update mode is a mapping $\mu$ which assigns, for any $n$ and any BN $f\in\mathrm{F}(n)$, a collection of trajectories closed by taking prefixes (i.e. if $(x^{0},\dots,x^{l})\in\mu(f)$, then $(x^{0},\dots,x^{a})\in\mu(f)$). Since we shall fix the BN in the remainder of this section, we shall omit the dependence on $f$ in our notation. To emphasize that $(x^{0},\dots,x^{l})$ is a trajectory according to $\mu$, we write $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}$. A pair of configurations $(x,y)$ is a reachability pair for an update mode $\mu$ if there exists a trajectory $x=x^{0}\to_{\mu}\dots\to_{\mu}x^{l}=y$, and we say that $y$ is reachable from $x$.
We aim to model updates where there are delays in the propagation of the updated values. Any trajectory can be represented using the starting configuration $x^{0}$ and for all $1\leq a\leq l$, a triple
$w^{a}=(i^{a},s^{a},t^{a})$
so that the next configuration is
$x^{a}=(f_{i^{a}}(s^{a}),t^{a}).$
Here is the intuition behind this representation and the kind of restrictions we shall apply to the word $w^{a}$. The component $i^{a}$ updating its state at step $a$ does not necessarily apply its update function $f_{i^{a}}$ to the current configuration $x^{a-1}$ but to a configuration $s^{a}$ built up of states with different timestamps. More precisely, we have $s^{a}$ with $s^{a}_{j}=x_{j}^{b_{j}}$ for all $j$, i.e. $s^{a}\in[x^{0},\dots,x^{a-1}]$. In the following, we will also allow for the other components to choose any value that has already been observed in the trajectory; we obtain a target configuration $t^{a}\in[x^{0},\dots,x^{a-1}]$.
Different update modes correspond to different constraints over $i^{a}$, $s^{a}$, and $t^{a}$. We list below the update modes that we shall consider in this paper. None of those applies any restriction on $i^{a}$: it can be any coordinate in $[n]$ at any time step. Note that in a general update mode $\mu$, we may have two consecutive trajectories $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}$ and $x^{l}=y^{0}\to_{\mu}\dots\to_{\mu}y^{m}$, but not their concatenation $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}=y^{0}\to_{\mu}\dots\to_{\mu}y^{m}$. This happens for instance if the source $s^{a}$ and the target $t^{a}$ are constrained to be equal to the starting configuration of the trajectory. This kind of situation seems rather artificial; as such, all the update modes that we list below are closed under concatenation of trajectories.
2.1 Asynchronous updates

### v2

@@
Abstract. Boolean networks are extensively applied as models of complex dynamical systems, aiming at capturing essential features related to causality and synchronicity of the state changes of components along time. Dynamics of Boolean networks result from the application of their Boolean map according to a so-called update mode, specifying the possible transitions between network configurations. In this paper, we explore update modes that possess a memory on past configurations, and provide a generic framework to define them. We show that recently introduced modes such as the most permissive and interval modes can be naturally expressed in this framework, and we propose novel update modes, the history-based, trapping, and subcube-based modes. Building on the unified definitions, we provide a comprehensive comparison of memory-based update modes, resulting in their hierarchy by simulation and weak simulation. Finally, we highlight consequences of introducing memory on the notions of trajectory and attractors.
@@ 1 Introduction
A large amount of theoretical work underlined the importance of the scheduling of the update of states of components for the dynamics that can be generated, the so-called update mode. One natural mode is the parallel (or synchronous) mode, where all components are updated simultaneously and instantaneously, generating deterministic dynamics. Other classical update modes are the (fully) asynchronous mode where only one component can be updated at a time, and the general asynchronous mode where any subset of components can be updated simultaneously, potentially leading to non-deterministic dynamics. By definition, the general asynchronous trajectories subsume the asynchronous and synchronous trajectories. The converse is, in general, false. Update modes also impact the so-called limit configurations, that are the configurations from which any reachable configuration can return to them. Limit configurations constitute the attractors of dynamics as they represent a set of configurations that, once reached, can no longer be evaded. Limit configurations are a prime subject of study in the literature, notably for determining their number and size, e.g., [10, 5, 6, 14], as well as their robustness to update modes, e.g., [2, 1].
When employing BNs as models of complex dynamical systems, such as gene regulation within cells, update modes aim at reflecting hypotheses related to the duration, speed, probability, and other quantitative features of the transitions. For instance, asynchronous modes can be motivated by the fact that genes can require different amounts of proteins to become active [20, 19].
@@
To that end, we propose a unifying framework for defining update modes with memory by characterizing the trajectories they generate. This allows to account for a sequence of configurations that have been computed since an initial one, and compute elongation of the trajectory based on this sequence. Besides reformulating above-mentioned update modes, we take advantage of the framework to introduce novel elementary update modes using memory: the history-based, the trapping, and the subcube-based update modes. Intuitively, the history-based mode extends the asynchronous by allowing to take any configuration of the past, instead of the latter only, to compute the next state of the component. Then, the trapping mode extends the history-based by allowing to take any configuration of the past as the base next configuration. Informally, this could be seen as some sort of time traveling: after the update of some components, one can return to a past configuration and enable state changes based on knowledge of the future (which can then be different). Finally, the subcube-based mode relies on the smallest subcube enclosing the configurations of the trajectory: the next configuration results from the update of any configuration of that subcube, with the state of one component computed from to any other configuration of the subcube.
@@
One can naturally derive that the asynchronous trajectories are a subset of history-based trajectories, being themselves a subset of trapping trajectories. In this paper, we establish simulation and weak simulation relationships with the interval, cuttable, most permissive and introduced update modes (Sect. 3; Fig. 2). In particular, we demonstrate that cuttable and history-based modes are incomparable, while both weakly simulating the interval, and both being weakly simulated by the most permissive, which in turns, is weakly simulated by the trapping mode. We also show that the trapping mode is weakly bisimilar to the subcube-based mode. We provide a similar hierarchy by considering inclusion of trajectories (simulation).
@@
Another important result is that most permissive and trapping modes have the same limit configurations, and have the same limit configurations reachable from any configuration (Sect. 4).
@@
The paper is structured as follows. Sect. 2 introduces our generic framework to express memory-based update modes which is employed to define novel trapping, history-based, and subcube-based update modes, and give equivalent definitions of asynchronous, most permissive, interval and cuttable update modes. Then, Sect. 3 provides a simulation and weak-simulation hierarchy between those updates, and Sect. 4 highlight important consequences of using memory in update modes on trajectories and attractors. Finally, Sect. 5 discusses the contributions of the paper.
@@
We denote the Boolean set by $\mathbb{B}=\{0,1\}$ and for any positive integer $n$, we denote $n]=\{1,\dots,n\}$. A configuration is $x=(x_{1},\dots,x_{n})\in\mathbb{B}^{n}$. For any $S\subseteq[n]$, we denote $x_{S}=(x_{s}:s\in S)$ and $x_{-S}=(x_{t}:t\notin S)$, and we use the notation $x=(x_{S},x_{-S})$. We shall identify an element $i\in[n]$ with the corresponding singleton $\{i\}$, so that $x=(x_{i},x_{-i})$ for instance. For any two configurations $x,y\in\mathbb{B}^{n}$, we denote the set of components where they differ by $\Delta(x,y)=\{i\in[n]:x_{i}\neq y_{i}\}$ and their Hamming distance by $d_{\mathrm{H}}(x,y)=|\Delta(x,y)|$. For any Boolean variable $a\in\mathbb{B}$, we denote its negation by $\neg a=1-a$; we extend this notation to configurations of any length by componentwise negation: $\neg x=(\neg x_{1},\dots,\neg x_{n})$.
@@
A subcube of $\mathbb{B}^{n}$ is any $X\subseteq\mathbb{B}^{n}$ such that there exist two disjoint sets of configurations $S,T\subseteq[n]$ with $X=\{x\in\mathbb{B}^{n}:x_{S}=0,x_{T}=1\}$. For any set $A\subseteq\mathbb{B}^{n}$, the principal subcube of $A$, denoted by $[A]$, is the smallest subcube containing $A$. If $A=\{a_{1},\dots,a_{k}\}$, we also denote $A]=[a_{1},\dots,a_{k}]$. If $X$ is a subcube and $x\in X$, then there is a unique $y\in X$ such that $X=[x,y]$; we refer to $y$ as the opposite of $x$ in $X$, and we denote it by y=Xxy=X-x.
@@
A Boolean network (BN) with $n$ components is a mapping $f:\mathbb{B}^{n}\to\mathbb{B}^{n}$. We denote the set of BNs with $n$ components as $\mathrm{F}(n)$. Any BN $f\in\mathrm{F}(n)$ can be viewed as $f=(f_{1},\dots,f_{n})$ where $f_{i}:\mathbb{B}^{n}\to\mathbb{B}$ is given by $f_{i}(x)=f(x)_{i}$ for all $i\in[n]$.
@@
A trapspace of $f\in\mathrm{F}(n)$ is a subcube $X\subseteq\mathbb{B}^{n}$ such that $f(X)\subseteq X$. The collection of all trapspaces of $f$ is closed under intersection. Then for any $x\in\mathbb{B}^{n}$, there is a smallest trapspace of $f$ that contains $x$, which we shall refer to as the principal trapspace of $x$ (with respect to $f$). For the sake of simplicity, we denote it by $T_{f}(x)$. The principal trapspace $T_{f}(x)$ can be recursively computed as follows: let $T_{0}=\{x\}$ and $T_{i}=[T_{i-1}\cup f(T_{i-1})]$, then $T_{n}=T_{f}(x)$. In particular, $x,f(x)]\subseteq T_{f}(x)$. A trapspace $T$ is minimal if there is no trapspace $T^{\prime}$ with $T^{\prime}\subsetneq T$. Clearly, any minimal trapspace is principal, but the converse does not necessarily hold.
@@
A (directed) graph is $\Gamma=(V,E)$, where $V$ is the set of vertices and $E\subseteq V^{2}$ is the set of edges. The asynchronous graph of a BN $f\in\mathrm{F}(n)$ is the graph $\mathtt{A}(f)=(V,E)$ where $V=\mathbb{B}^{n}$ and
$E=\{(f_{i}(x),x_{-i}):x\in\mathbb{B}^{n},i\in[n]\}.$
@@
Note that in most literature, one removes the loops $(x,x)$ from the asynchronous graph, that occur every time $f_{i}(x)=x_{i}$, yet we shall keep those loops instead in our definition. However, when drawing the asynchronous graph, we shall not display the loops and instead draw the underlying hypercube with thin black lines and the arcs of the graph with thick blue arrows. See below for an example of a BN, for which we give the asynchronous graph and all the principal and minimal trapspaces.
@@
000000001001010010011011100100101101110110111111
Then the principal trapspaces of $f$ are given below, where $T_{f}(100)$, $T_{f}(101)$ and $T_{f}(110)$ are the only minimal trapspaces of $f$:
$T_{f}(000)$ $=[000,110]$ $=\{x:x_{3}=0\}$
$T_{f}(001)$ $=[001,110]$ $=\{x\}=\mathbb{B}^{3}$
$T_{f}(010)$ $=[010,100]$ $=\{x:x_{3}=0\}$
$T_{f}(011)$ $=[011,100]$ $=\{x\}=\mathbb{B}^{3}$
$T_{f}(100)$ $=[100,100]$ $=\{100\}$
$T_{f}(101)$ $=[101,101]$ $=\{101\}$
$T_{f}(110)$ $=[110,110]$ $=\{110\}$
$T_{f}(111)$ $=[111,110]$ $=\{x:x_{1}=x_{2}=1\}.$
@@ 2.1 A unified framework
Before giving the unified framework in its generality, we would like to illustrate it via the example of an asynchronous trajectory. Consider the network $f$ from Example 1.1. By following the arcs of the asynchronous graph, we have a trajectory
$x^{0}=011\to x^{1}=010\to x^{2}=000\to x^{3}=100\to x^{4}=100.$
Each update is of the form
$x^{a-1}=(x^{a-1}_{i^{a}},x^{a-1}_{-i^{a}})\to x^{a}=(f_{i^{a}}(x^{a-1}),x_{-i^{a}}),$
where $i^{a}$ is the component updated at time $a$. We note that $i^{4}$ could be any index in $[n]$, since $x^{3}$ is a fixed point, but this is inconsequential. As such, we can succinctly represent each update by the index $i^{a}\in[n]$, and the whole trajectory can be obtained from its starting configuration $x^{0}=011$ and the sequence $i^{1}=3,i^{2}=2,i^{3}=1,i^{4}=1$.
In order to add memory to our model of trajectories, we will allow two generalisations:
  1. 1.
    the function $f_{i^{a}}$ may be applied to a source configuration $s^{a}\in\mathbb{B}^{n}$, which may be distinct from $x^{a-1}$, so that $x^{a}_{i^{a}}=f_{i^{a}}(s^{a})$;
  2. 2.
    the rest of the configuration can be taken from a target configuration $t^{a}\in\mathbb{B}^{n}$, which again may be distinct from $x^{a-1}$, so that $x^{a}_{-i^{a}}=t^{a}_{-i^{a}}$
As such, the transition becomes
$x^{a-1}=(x^{a-1}_{i^{a}},x^{a-1}_{-i^{a}})\to x^{a}=(f_{i^{a}}(s^{a}),t^{a}_{-i^{a}}),$ (1)
and it can be represented by the triple
$w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}.$
Again, different choices for $w^{a}$ may lead to the same next configuration $x^{a}$, but an entire trajectory $x^{0}\to\dots\to x^{l}$ can be obtained from $x^{0}$ and the sequence of triples $w^{1},\dots,w^{l}$. For instance, the asynchronous trajectory above can be obtained from its starting configuration $x^{0}=011$ and the sequence of triples
$w^{1}=(3,011,011),\quad w^{2}=(2,010,010),\quad w^{3}=(1,000,000),\quad w^{4}=(1,100,100).$
@@
In its full generality, a (fully asynchronous) trajectory is a sequence $(x^{0},\dots,x^{l})\in(\mathbb{B}^{n})^{*}$ such that for all $1\leq a\leq l$, there exists $t\in[x^{0},\dots,x^{a-1}]$ such that $d_{\mathrm{H}}(t,x^{a})\leq 1$. In particular, a geodesic is a trajectory $(x^{0},\dots,x^{l})$ such that $d_{\mathrm{H}}(x^{a-1},x^{a})=1$ for all $1\leq a\leq l$ and $l=d_{\mathrm{H}}(x^{0},x^{l})$; this is equivalent to a shortest path from $x^{0}$ to $x^{l}$ in the hypercube. A prefix of the trajectory $(x^{0},\dots,x^{l})$ is any $(x^{0},\dots,x^{a})$ for $0\leq a\leq l$, or the empty sequence. An update mode is a mapping $\mu$ which assigns, for any $n$ and any BN $f\in\mathrm{F}(n)$, a collection of trajectories closed by taking prefixes (i.e. if $(x^{0},\dots,x^{l})\in\mu(f)$, then $(x^{0},\dots,x^{a})\in\mu(f)$). Since we shall fix the BN in the remainder of this section, we shall omit the dependence on $f$ in our notation. To emphasize that $(x^{0},\dots,x^{l})$ is a $\mu$-trajectory, we write $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}$. A pair of configurations $(x,y)$ is a reachability pair for an update mode $\mu$ if there exists a $\mu$-trajectory $x=x^{0}\to_{\mu}\dots\to_{\mu}x^{l}=y$, and we say that $y$ is reachable from $x$ and denote it with $x\to^{*}_{\mu}y$.
In this paper, we consider update modes that satisfy certain constraints. In particular, we assume that $\mu$ only produces trajectories where each update is of the form of Equation (1) for some reasonable choices of source and target configurations. More precisely, the trajectories are $(x^{0},\dots,x^{l})$, where for each $1\leq a\leq l$, there exists $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ such that:
  1. 1.
    $s^{a}\in[x^{0},\dots,x^{a-1}]$ is a configuration made up of states with different time stamps from the trajectory, so that $f_{i^{a}}$ is applied to state values that are drawn from the trajectory’s history;
  2. 2.
    similarly, $t^{a}\in[x^{0},\dots,x^{a-1}]$, so that the remainder of the $x^{a}$ configuration is made up of state values drawn from the trajectory’s history;
  3. 3.
    the configuration $x^{a}$ is given by $x^{a}=(f_{i^{a}}(s^{a}),t^{a}_{-i^{a}})$.
Different update modes correspond to different constraints over $i^{a}$, $s^{a}$, and $t^{a}$ (the case without any further constraints will be the so-called subcube update mode). We list below the update modes that we shall consider in this paper. None of those applies any restriction on $i^{a}$: it can be any component in $[n]$ at any time step. We shall highlight some important properties of these update modes with memory in Section 4, once they have been presented and thoroughly compared.
2.2 Update modes
2.2.1 Asynchronous updates
We begin with the simplest update mode, namely asynchronous updates [20]. Let $f\in\mathrm{F}(n)$. The asynchronous updating mode considers $x\to y$ iff $\exists i\in[n]$ such that $y_{i}=f_{i}(x)$ and $y_{-i}=x_{-i}$.